What would the future value of $6.5 million dollars be after 5 years at 1.2% interest rate?
1
Expert's answer
2011-05-17T09:27:47-0400
FV = PV (1 +r)t
FV = 6.5 (1 + 0.012)5 = $6.9 million dollars